# A Quality Diversity Method to Automatically Generate Multi-Agent Path Finding Benchmark Maps

### v4

@@ Abstract
We use the Quality Diversity (QD) algorithm with Neural Cellular Automata (NCA) to generate benchmark maps for Multi-Agent Path Finding (MAPF) algorithms. Previously, MAPF algorithms are tested using fixed, human-designed benchmark maps. However, such fixed benchmark maps have several problems. First, these maps may not cover all the potential failure scenarios for the algorithms. Second, when comparing different algorithms, fixed benchmark maps may introduce bias leading to unfair comparisons between algorithms. Third, since researchers test new algorithms on a small set of fixed benchmark maps, the design of the algorithms may overfit to the small set of maps. In this work, we take advantage of the QD algorithm to (1) generate maps with patterns to comprehensively understand the performance of MAPF algorithms, (2) be able to make fair comparisons between two MAPF algorithms, providing further information on the selection between two algorithms and on the design of the algorithms. Empirically, we employ this technique to generate diverse benchmark maps to evaluate and compare the behavior of different types of MAPF algorithms, including search-based, priority-based, rule-based, and learning-based algorithms. Through both single-algorithm experiments and comparisons between algorithms, we identify patterns where each algorithm excels and detect disparities in runtime or success rates between different algorithms.
1 Introduction
@@
We study the problem of generating diverse and targeted benchmark maps for Multi-Agent Path Finding (MAPF) algorithms. Given a map and a group of agents, MAPF is the problem of finding collision-free paths from their start to goal locations. MAPF has wide applications in coordinating hundreds of robots in automated warehouses Li et al. (2021c); Zhang et al. (2023a, b), moving characters in video games Ma et al. (2017), managing drone traffic Choudhury et al. (2022), and controlling multi-robotic arms Shaoul et al. (2024).
@@
Given the wide applicability of MAPF, many algorithms are proposed to solve MAPF, and benchmarking these algorithms becomes an important task. Stern et al. (2019) have proposed a set of 33 MAPF benchmark maps, covering a diverse spectrum of map sizes, layouts, and difficulties [^1] https://movingai.com/benchmarks/mapf/index.html. These human-designed maps are widely used to test newly proposed MAPF algorithms Li et al. (2021b, a); Okumura et al. (2019); Skrynnik et al. (2024). However, such benchmark maps have several problems. First, they may not fully encompass the failure modes of an algorithm. Second, with a small fixed set of maps, it is challenging to sufficiently understand the pros and cons of different algorithms in different maps. Third, while comparing multiple algorithms, researchers are prone to run experiments on a subset of maps that are biased towards the proposed algorithms, resulting in an unfair comparison.
@@
Meanwhile, Quality Diversity (QD) algorithms have been used to generate a diverse set of high-quality solutions by optimizing a given objective function and a set of diversity measure functions. A recent work Zhang et al. (2023b) has used QD algorithms to optimize layouts for automated warehouses. To ensure that the optimized layouts possess human-explainable and regularized patterns, a follow-up work Zhang et al. (2023a) then leverages Neural Cellular Automata (NCA) to generate the layouts, while applying QD algorithms to optimize the parameters of the NCA. Cellular Automata Gardner (1970) iteratively generates complex tile-based structures from a simple one through local interaction between tiles. Each tile decides its next state based on its neighbors via a fixed rule. Neural Cellular Automata (NCA) represents the rule using a convolutional neural network.
In this paper, we adapt the layout optimization approach based on the QD algorithm and NCA from the previous work Zhang et al. (2023a) and use it with an alternative goal of generating diverse benchmark maps for MAPF algorithms. To demonstrate that our approach can generate benchmark maps for a broad spectrum of MAPF algorithms, we choose five algorithms, namely CBS Sharon et al. (2015), EECBS Li et al. (2021b), PBS Ma et al. (2019), PIBT Okumura et al. (2019), and Learn-to-Follow (LTF) Skrynnik et al. (2024), representing search-based, priority-based, rule-based, and learning-based algorithms, respectively.
We first generate benchmark maps for each algorithm individually, presenting diverse maps that are easy or challenging for each algorithm to solve. We summarize the design principles of benchmark maps that target each algorithm. We then generate benchmark maps that automatically compare two MAPF algorithms by generating maps that are easy for one algorithm and hard for the other. We aim to understand in which maps one algorithm outperforms the other in terms of either runtime or success rate.
@@
We make the following contributions: (1) we adapt the layout optimization approach based on QD algorithms, proposing a framework to generate diverse MAPF benchmark maps, (2) we apply our framework on five representative MAPF algorithms of different categories, generating novel benchmark maps and providing new insights about their performance on maps of different patterns, and (3) we give guidelines for future researchers on how to better benchmark MAPF algorithms.
2 Preliminaries
2.1 Multi-Agent Path Finding (MAPF)
@@
Given a valid map and a set of agents with their start and goal locations, MAPF aims to find collision-free paths from their start to goal locations. Agents can either move to their adjacent vertices or stay at their current vertices at each timestep. Two agents collide if they are at the same vertex or swap vertices at the same timestep. The objective of MAPF is minimizing the sum-of-cost, defined as the sum of path lengths of all agents.
@@
Search-Based. This category includes exponential-time algorithms that exhaustively explore the solution space of MAPF. They usually have theoretical guarantees such as optimality and suboptimality, but suffer from computational time. Examples include M Wagner and Choset (2015), BCP Lam et al. (2019), CBS Sharon et al. (2015); Li et al. (2020), ECBS Barer et al. (2014) and EECBS Li et al. (2021b). We choose CBS and EECBS as the representatives of this category. CBS starts by planning a shortest path for each agent, ignoring the collisions, and then resolving the collisions with a two-level search. The high-level search iteratively selects unresolved collisions and adds constraints to tackle them. The low-level search runs single-agent planning to compute new paths that satisfy the constraints. The search continues until all collisions are resolved. EECBS is a suboptimal variant of CBS that uses Explicit Estimation Search Thayer and Ruml (2011) on the high level and focal search Pearl and Kim (1982) on the low level.
@@
Priority-Based. Priority-based algorithms plan paths for each agent in a sequence, forcing agents with low priority to avoid colliding with those with high priority. The priority planning (PP) Erdmann and Lozano-Pérez (1987) algorithm plans paths for agents in a pre-defined priority order. Monte-Carlo PP Friedrich et al. (2024) pre-defines a set of randomized priority orders and selects the paths with the best solutions. PBS Ma et al. (2019) combines CBS with PP to explicitly search for a good priority order with a two-level search. The high-level search of PBS is similar to CBS except that PBS constrains one agent to have a higher priority order than the other. The low-level search then plans paths with the order from the high-level search. We choose PBS as the representative of this category.
@@
Rule-Based. Rule-based algorithms leverage pre-defined rules to move agents to their goals. They usually run much faster but produce worse solutions than search-based and priority-based methods. We choose PIBT Okumura et al. (2019) as the representative of this category. PIBT uses an iterative one-timestep rule to move the agents. At each timestep, each agent plans a single-step action towards its goal. In case of collision, a pre-defined rule based on priority inheritance and backtracking is applied to resolve the collision.
@@
Learning-Based. Learning-based algorithms usually formulate MAPF as a Multi-Agent Reinforcement Learning (MARL) Sartoretti et al. (2019) problem. They train a shared policy for each agent, taking its local field of view as input and deciding its next action. We choose Learn to Follow (LTF) Skrynnik et al. (2024) as the representative of this category. LTF starts by searching for a guide path for each agent without considering collisions. Then it uses a shared learnable policy to move the agents to their goals along the guide paths while avoiding collisions. LTF is developed for lifelong MAPF, a variant of MAPF that constantly assigns new goals to agents. We modify it for MAPF by asking agents to stop when they reach their first goals.
2.2 MAPF Benchmarks
Early works in MAPF typically use randomly generated benchmark maps Wagner and Choset (2015) or maps from the single-agent path planning benchmark set Sturtevant (2012). Stern et al. (2019) established the major set of MAPF benchmark maps with 33 maps, which is used extensively in MAPF research Li et al. (2021a); Okumura (2023); Friedrich et al. (2024). However, researchers tend to select a subset of all the maps while testing novel MAPF algorithms. This could result in unfair and biased benchmarking results because they can intentionally or unintentionally cherry-pick maps that favor their proposed algorithms, or the design of their algorithms could overfit to the selected maps. The MARL community usually generates random maps to test the trained policy for MAPF problems Sartoretti et al. (2019). Researchers who develop MAPF algorithms for the multi-robot systems for automated warehouses or sortation centers often use a set of warehouse maps Li et al. (2021c); Zhang et al. (2023b, a, 2024).
2.3 Quality Diversity (QD) Algorithms
QD algorithms Cully et al. (2015) are inspired by evolutionary algorithms with diversity search to generate a diverse collection of high-quality solutions by optimizing an objective function and diversifying a set of diversity measure functions. QD algorithms maintain an archive, which is a tessellated measure space defined by the measure functions. The archive stores the best solution in each tessellated cell. QD algorithms then optimize the sum of objective values of all solutions in the archive, defined as QD-score. We choose Covariance Matrix Adaptation MAP-Annealing (CMA-MAE) Fontaine and Nikolaidis (2023) as the method for generating benchmark maps because it is the state-of-the-art QD algorithm specialized for continuous search domains. CMA-MAE is an extension of MAP-Elites Mouret and Clune (2015) that incorporates the covariance matrix adaptation mechanism of CMA-ES Hansen (2016), which is a derivative-free single-objective optimizer. CMA-ES maintains a multi-variate Gaussian distribution and iteratively samples from it for new solutions. It evaluates the solutions and updates the Gaussian toward high-objective regions. CMA-MAE adapts this mechanism to optimize the QD-score.
2.4 Automatic Scenario Generation
Automated methods of generating scenarios to understand, test, or benchmark a given algorithm have been proposed by different fields. In autonomous driving, researchers have generated scenarios to test and evaluate developed autonomous driving systems Abeysirigoonawardena et al. (2019). In human-robot-interaction, prior works have used QD algorithms to generate shared autonomy scenarios Fontaine et al. (2021b) or diverse kitchen layouts Fontaine et al. (2021a) to understand the coordination behavior between humans and robots. A follow-up work Bhatt et al. (2023) leverages model-based QD methods Bhatt et al. (2022); Zhang et al. (2022) to improve the sample efficiency of QD algorithms. In reinforcement learning, prior works have generated maps to benchmark Cobbe et al. (2020) or continuously improve Wang et al. (2019) trained agents.
@@ 2.4 Automatic Scenario Generation
In MAPF, one recent work Ren et al. (2024) has used QD to generate maps for MAPF. Our work differs in two key aspects: First, they focus on generating maps of fixed difficulties, quantified by an approximate metric based on map connectivity, while we are interested in generating benchmarking maps targeted for specific MAPF algorithms. Second, they directly optimize tile types (i.e., whether it is an empty space or an obstacle) using MAP-Elites, while we optimize a map generator based on NCA, which has been shown to be more effective for generating maps with diverse regularized patterns Zhang et al. (2023a).
3 Benchmark Generation Approach
Figure 1: Overview of our approach of using CMA-MAE to optimize NCAs to generate diverse benchmark maps.
@@ 3 Benchmark Generation Approach
We adapt the previous work Zhang et al. (2023a) to use CMA-MAE and NCA to generate diverse benchmark maps with the objective and measures computed by running MAPF algorithms. Figure 1 shows the overview of the method. Starting by sampling $b$ parameter vectors $\boldsymbol{\theta}$, we form $b$ NCA generators and use them to generate $b$ maps. Maps generated by NCA might not be valid. We then adapt a mixed integer linear programming (MILP) solver Zhang et al. (2023b) to repair the map such that (1) the map is valid, and (2) the number of obstacles falls in a pre-defined range $[O_{lb},O_{ub}]$. We then run MAPF algorithms to evaluate the objective and measure values and add the evaluated maps to an archive. For each evaluation, we run the given MAPF algorithm in $N_{e}$ instances. Finally, we update the Gaussian distribution, starting a new iteration. We repeat this process until we evaluate $N_{eval}$ maps.
@@
MAPF Instance. In this paper, we focus on generating MAPF benchmark maps of a pre-defined size and number of agents. To generate a MAPF instance based on a map, we follow the bucket method Stern et al. (2019) to generate evenly distributed start and goal locations of the agents with a distance constraint, which was used for generating the even scenarios in the MAPF benchmark. For each map, we generate $N_{e}$ MAPF instances with different starts and goals, run MAPF algorithms on each instance, and compute the average objective and measures of all instances as the objective and measures of the map.
@@
Framework Realization. Our proposed framework is flexible on a high level. It is capable of generating a diverse set of high-quality MAPF benchmark maps tailored to different MAPF algorithms by carefully designing objective and diversity measures. In this section, we propose two concrete realizations of our framework that are useful for benchmarking MAPF algorithms by designing objectives and measures. We first provide an one-algorithm realization, intending to generate diverse and difficult benchmark maps targeted for five representative MAPF algorithms. We then provide a two-algorithm realization, aiming to generate maps that are easy for one algorithm and hard for the other. We highlight that other realizations, such as generating easy maps or comparing more than two algorithms, are possible.
3.1 Objective

### v5

# QD-MAPPER: A Quality Diversity Framework to Automatically Evaluate Multi-Agent Path Finding Algorithms in Diverse Maps

@@ Abstract.
We use the Quality Diversity (QD) algorithm with Neural Cellular Automata (NCA) to automatically evaluate Multi-Agent Path Finding (MAPF) algorithms by generating diverse maps. Previously, researchers typically evaluate MAPF algorithms on a set of specific, human-designed maps at their initial stage of algorithm design. However, such fixed maps may not cover all scenarios, and algorithms may overfit to the small set of maps. To seek further improvements, systematic evaluations on a diverse suite of maps are needed. In this work, we propose Quality-Diversity Multi-Agent Path Finding Performance EvaluatoR (QD-MAPPER), a general framework that takes advantage of the QD algorithm to comprehensively understand the performance of MAPF algorithms by generating maps with patterns, be able to make fair comparisons between two MAPF algorithms, providing further information on the selection between two algorithms and on the design of the algorithms. Empirically, we employ this technique to evaluate and compare the behavior of different types of MAPF algorithms, including search-based, priority-based, rule-based, and learning-based algorithms. Through both single-algorithm experiments and comparisons between algorithms, researchers can identify patterns that each MAPF algorithm excels and detect disparities in runtime or success rates between different algorithms. Our generated maps are available at https://airtclick.github.io/qdmapper. Our code is available at https://github.com/AirTClick/QD-MAPPER.
Key words and phrases:
Multi-Agent Path Finding; Procedural Content Generation; Quality Diversity Optimization
1. Introduction
@@
We study the problem of evaluating the performance of Multi-Agent Path Finding (MAPF) algorithms by generating diverse maps. Given a map and a group of agents, MAPF is the problem of finding collision-free paths from their start to goal locations. MAPF has wide applications in coordinating hundreds of robots in automated warehouses ZhangNCA2023; zhangLayout23, moving characters in video games MaAIIDE17, managing drone traffic choudhury2022TruckUAVmapf, and controlling multi-robotic arms ShaoulICAPS24.
@@
Given the wide applicability of MAPF, many algorithms have been proposed to solve MAPF, and evaluating these algorithms becomes an important task. At the beginning of the algorithm design, researchers prefer to use benchmark maps to have an overview on the performance of their algorithms. SternSoCS19 have proposed a set of 33 MAPF benchmark maps, covering a diverse spectrum of map sizes, layouts, and difficulties. These human-designed maps are widely used to test newly proposed MAPF algorithms Li2020EECBSAB; okumura2019priority; learntofollow2024. However, to further discover the pros and cons and improve algorithms, such fixed maps are not sufficient.
@@
Meanwhile, Quality Diversity (QD) algorithms have been used to generate a diverse set of high-quality solutions by optimizing a given objective function and a set of diversity measure functions. A recent work zhangLayout23 has used QD algorithms to optimize layouts for automated warehouses. To ensure that the optimized layouts possess human-explainable and regularized patterns, a follow-up work ZhangNCA2023 then leverages Neural Cellular Automata (NCA), a CNN-based neural generator, to generate the layouts and uses QD algorithms to optimize the parameters of the NCA. Cellular Automata gardner1970 iteratively generates complex cell-based structures from a simple one through local interaction between cells. Each cell decides its next state based on its neighbors via a fixed rule. NCA then represents the rules using a convolutional neural network that can be optimized.
In this paper, we adapt the layout optimization approach from the previous work ZhangNCA2023, proposing Quality-Diversity Multi-Agent Path Finding Performance EvaluatoR (QD-MAPPER), a general framework with the goal of evaluating MAPF algorithms by generating diverse maps. To demonstrate that our approach can evaluate a broad spectrum of MAPF algorithms, we present evaluation results on 6 representative MAPF algorithms, namely CBS SharonAIJ15, EECBS Li2020EECBSAB, PBS MaAAAI19, LaCAM3 okumura2024lacam3, PIBT okumura2019priority, and Learn-to-Follow (LTF) learntofollow2024. We present a number of interesting results undiscovered in prior works. For example, we generate new maps with patterns that LaCAM3 runs out of time, while it was reported to solve 99% of the MAPF benchmarks SternSoCS19. We also expose the incompleteness of PBS by presenting maps in which PBS fails to find any solution instead of running out of time on benchmark maps.
@@
We make the following contributions: (1) we adapt the layout optimization approach to propose QD-MAPPER to evaluate MAPF algorithms by generating diverse maps, (2) we apply QD-MAPPER on 6 representative MAPF algorithms of different categories, finding new challenging maps for MAPF algorithms, and providing new insights about their performance on maps of different patterns, and (3) we propose two concrete realizations of QD-MAPPER that are useful for evaluating MAPF algorithms: one-algorithm realization and two-algorithm realization. As a generic framework, researchers can extend other realizations, such as generating easy maps or comparing more than two algorithms.
2. Preliminaries
2.1. Multi-Agent Path Finding (MAPF)
@@
Given a valid map and a set of agents with their start and goal locations, MAPF aims to find collision-free paths from their start to goal locations. Agents can either move to their adjacent locations or stay at their current locations at each timestep. Two agents collide if they are at the same location or swap locations at the same timestep. The objective of MAPF is minimizing the sum-of-cost, defined as the sum of travel time of all agents.
@@
Search-Based. This category includes exponential-time algorithms that exhaustively explore the solution space of MAPF. They usually have theoretical guarantees such as optimality and bounded suboptimality, but suffer from long computational times. Examples include M MStar, BCP LamIJCAI19, ICTS SharonAIJ13, CBS SharonAIJ15, ECBS BarrerSoCS14, and EECBS Li2020EECBSAB . We choose CBS SharonAIJ15 and EECBS Li2020EECBSAB as the representatives of this category. CBS is an optimal algorithm. It starts by planning a shortest path for each agent, ignoring the collisions, and then resolving the collisions with a two-level search. The high-level search iteratively selects unresolved collisions and adds constraints to tackle them. The low-level search runs single-agent planning to compute new paths that satisfy the constraints. The search continues until all collisions are resolved. EECBS is a bounded-suboptimal variant of CBS that uses Explicit Estimation Search Thayer2011EES on the high level and focal search PearTPAMIl1982 on the low level.
@@
Priority-Based. Priority-based algorithms plan paths for each agent following a priority order, forcing agents with lower priorities to avoid colliding with those with higher ones. The priority planning (PP) Erdmann87; StandleyAAAI10 algorithm plans paths for agents in a pre-defined priority order. Monte-Carlo PP friedrich2024scalable pre-defines a set of randomized priority orders and selects the one with the best paths. Representative algorithms include priority planning (PP) Erdmann87, Monte-Carlo PP friedrich2024scalable, and PBS MaAAAI19. We choose PBS as the representative of this category. PBS combines CBS with PP to explicitly search for a good priority order with a two-level search. The high-level search of PBS is similar to CBS except that PBS constrains one agent to have a higher priority order than the other. The low-level search then plans paths with the order from the high-level search.
@@
Rule-Based. Rule-based algorithms WangB11 leverage pre-defined rules to move agents to their goals. They usually run much faster, but produce worse solutions than search-based and priority-based methods. We choose PIBT okumura2019priority and LaCAM3 okumura2024lacam3 as representatives. PIBT uses an iterative one-timestep rule to move agents. At each timestep, each agent plans a single-step action towards its goal. In case of collisions, a pre-defined rule is applied to resolve the collision. LaCAM3 is an anytime algorithm that combines PIBT and search-based methods. As a hybrid algorithm that combines PIBT with search-based methods, It performs the search by sequentially expanding search nodes, and when it encounters the goal configuration, it derives the solution by backtracking parent pointers.
@@
Learning-Based. Learning-based algorithms either formulate MAPF as a Multi-Agent Reinforcement Learning (MARL) guillaume2018primal problem or leverage Imitation Learning (IL) to imitate the behavior of an oracle MAPF algorithm. MARL-based methods guillaume2018primal; Wang2023SCRIMPSC; learntofollow2024 typically train a shared policy for each agent, taking its local field of view as input and deciding its next action. IL-based methods, on the other hand, attempt to imitate an oracle algorithm. For example, MAPF-GPT Andreychuk_MAPFGPT_2025 tokenizes MAPF instances and trains a transformer to imitate LaCAM3 okumura2024lacam3. SILLM JiangICRA25 trains a learnable PIBT by imitating the behavior of WPPL Jiang2024Competition. We choose Learn to Follow (LTF) learntofollow2024 as the representative of the learning-based algorithms. LTF starts by searching for a guide path for each agent without considering collisions. Then it uses a shared learned policy to move the agents to their goals along the guide paths while avoiding collisions. LTF is developed for lifelong MAPF, a variant of MAPF that constantly assigns new goals to agents. We modify it for MAPF by asking agents to stop when they reach their first goals.
2.2. MAPF Algorithm Evaluation
In early works, researchers typically used randomly generated benchmark maps MStar or maps from the single-agent path planning benchmark set sturtevant2012benchmarks to evaluate new MAPF algorithms. SternSoCS19 established the major set of MAPF benchmark maps with 33 maps, which are used extensively in MAPF research Li2020EECBSAB; Chan2023GreedyPS; zhang2024ggo. These maps can provide an overview of the performance of MAPF algorithms at the initial stage. However, to further discover the pros and cons of the algorithm, the benchmark set is insufficient with its fixed patterns and limited size. For example, LaCAM3, despite prior claims of solving 99% of standard benchmarks, also struggles under our generated maps with long corridors and narrow entries, which are patterns underrepresented in existing benchmarks.
@@ 2.2. MAPF Algorithm Evaluation
To the best of our knowledge, very few works use automatic generation methods to evaluate the performance of MAPF algorithms. The POGEMA platform skrynnik2025pogema leverages procedural content generation Cobbe2019LeveragingPG to train and evaluate MARL agents, but it does not target specific MAPF algorithms. One recent work ren2024MAPFhard also uses QD to generate maps for MAPF. However, it has two major differences with our work. First, they focus on generating maps of fixed difficulties, quantified by an approximate metric based on map connectivity, while we are interested in evaluating the performance of specific MAPF algorithms by generating diverse maps with different difficulties. Second, they directly optimize tile types (i.e., whether it is an empty space or an obstacle), while we optimize a map generator based on NCA, which has been shown to be more effective for generating maps with diverse regularized patterns ZhangNCA2023.
2.3. Quality Diversity (QD) Algorithms and Automatic Scenario Generation
QD algorithms mouret2015illuminating are inspired by evolutionary algorithms with diversity search to generate a diverse collection of high-quality solutions by optimizing an objective function and diversifying a set of diversity measure functions. QD algorithms maintain an archive, which is a tessellated measure space defined by the measure functions. The archive stores the best solution in each tessellated cell. QD algorithms then optimize the sum of objective values of all solutions in the archive, defined as QD-score. We choose Covariance Matrix Adaptation MAP-Annealing (CMA-MAE) Fontaine2022CovarianceMA as the method to generate maps because it is the state-of-the-art QD algorithm specialized for continuous search domains. CMA-MAE is an extension of MAP-Elites mouret2015illuminating that incorporates the covariance matrix adaptation mechanism of CMA-ES hansen2016cmaes, which is a derivative-free single-objective optimizer. CMA-ES maintains a multi-variate Gaussian distribution and iteratively samples from it for new solutions. It evaluates the solutions and updates the Gaussian toward high-objective regions. CMA-MAE adapts this mechanism to optimize the QD-score. QD algorithms are widely applied in automatic scenario generation scenarios. In multi-robot systems, ZhangNCA2023 optimize NCA via QD algorithms to generate arbitrarily large warehouse layouts. In autonomous driving, researchers generate scenarios to evaluate developed autonomous driving systems Abeysirigoonawardena2019GeneratingAD; Mullins2018AdaptiveGO. In human-robot interaction, QD algorithms are used to generate shared autonomy scenarios fontaine2021quality and diverse kitchen layouts fontaine2021importance to study the coordination behavior between humans and robots. A follow-up work Bhatt2023surrHRI leverages model-based QD methods Bhatt2022DeepSA; Zhang2021DeepSA to improve the sample efficiency of QD algorithms. In reinforcement learning, prior works have generated maps to benchmark Cobbe2019LeveragingPG or continuously improve Wang2019PairedOT trained agents.
3. Map Generation Approach
Figure 1. Overview of QD-MAPPER: (1) sample parameter vectors $\boldsymbol{\theta}$ from multi-variate Gaussian of CMA-MAE, (2) form NCA generators, (3) use a Mixed Integer Linear Programming (MILP) solver to ensure map connectivity and the correct number of obstacles, (4) run MAPF algorithms, (5) update the archive, and (6) update multi-variate Gaussian.
[…]
[Overview of QD-MAPPER]The pipeline shows how QD-MAPPER generates maps to evaluate the performance of targeted algorithms.
3.1. Overview
Figure 1 shows the pipeline of QD-MAPPER. We adapt previous works to use CMA-MAE Fontaine2022CovarianceMA to search for and update a diverse collection of NCA ZhangNCA2023 generators to generate diverse maps with the objective and measures computed by running MAPF algorithms. We start by sampling a batch of $b$ parameter vectors $\theta$ from a multivariate Gaussian distribution, forming $b$ NCA generators. Starting from a fixed seed map, each NCA generator iteratively updates the input map to generate a map with complex local patterns, leading to $b$ maps. Appendix A exemplifies the process of generating a map.
@@ 3.1. Overview
Maps generated by NCA might not be valid. We then adapt a Mixed Integer Linear Programming (MILP) solver zhang:aiide2020; zhangLayout23 to repair the map to enforce connectivity and domain-specific constraints while making minimal modifications to the generated map. We then evaluate repaired maps with MAPF algorithms to calculate the average objectives and measure values, and add the evaluated maps to an archive. For each evaluation, we run the given MAPF algorithm in $N_{e}$ instances. Finally, we update the parameters of the multivariate Gaussian distribution, sampling a new batch of $b$ parameter vectors and starting a new iteration. We repeat this process until we evaluate $N_{eval}$ maps.
NCA generator. We define an NCA generator as a function $g(s;\theta;C):M\to M$, where $M$ is the space of possible maps, $s\in M$ represents a fixed seed map, and $C\in\mathbb{Z}^{+}$ is a hyperparameter controlling the number of iterations to run the NCA generator. The generator is parameterized by $\theta\in\Theta$, where each parameter vector $\theta$ corresponds to a distinct NCA generator that then defines a distinct map $m\in M$. We then use CMA-MAE to optimize $\theta$, attempting to find the best NCA generator in each cell of the archive.
Our NCA generator is a convolutional neural network (CNN) following the architecture in prior work ZhangNCA2023. It consists of three convolutional layers with kernel size $3\times 3$, each followed by either ReLU or sigmoid activations. We represent the 2D tile-based maps as 3D tensors, with each tile encoded as a one-hot vector, representing its tile type (i.e. obstacle or empty space). Since the output of the NCA has the same size as its input, we can repeatedly pass the output back into the network.
MILP Solver. Maps generated by NCA might not be valid (e.g., a wall partitioning two open sections). To ensure that the maps are valid, we formulate and solve a MILP zhang:aiide2020; zhangLayout23 to obtain a repaired map. The MILP constraints are set to be: (1) all empty spaces are connected, and (2) the number of obstacles falls within a pre-defined range $O_{lb},O_{ub}]$. We set the objective to minimize the Hamming distance between unrepaired and repaired maps to ensure that the patterns of the generated map are maintained as much as possible.
@@
Generation of MAPF Instances From Maps. To generate a MAPF instance based on a map, we follow the bucket method SternSoCS19 to generate evenly distributed start and goal locations of the agents with a distance constraint, which was used for generating the even scenarios in the MAPF benchmark. For each map, we generate $N_{e}$ MAPF instances with different starts and goals, run MAPF algorithms on each instance, and compute the average objective and measures of all instances as the objective and measures of the map. Notably, our MAPF instance generation strategy serves only as an example. In applications where a different set of instances is needed, users can customize the instance generation method inside QD-MAPPER.
@@
Framework Realization. Our proposed framework is flexible at a high level. In this paper, we propose two concrete realizations of QD-MAPPER that are useful for evaluating MAPF algorithms with carefully designed objectives and measures. We first provide a one-algorithm realization, intending to evaluate five representative MAPF algorithms by generating diverse and difficult maps for each of them. We then provide a two-algorithm realization, aiming to compare the performance of two algorithms by generating maps that are easy for one algorithm and hard for the other. We highlight that other realizations, such as generating easy maps or comparing more than two algorithms, are possible.
3.2. Objective